OpenStudy (anonymous):

Find the exact value of each of the other five trig functions (besides the function tan) for the angle x - without finding x - given that the terminal side of x is in quadrant 2 an tan(x)=-3

OpenStudy (anonymous):

OK so this is how i started working this one out... and I think one of my first steps is wrong.

OpenStudy (anonymous):

tan= r/b ...therefore if given tan(x)=-3, then expressed in another way, tan=-3/1

OpenStudy (anonymous):

aww crap... i mean... tan=b/a ... but i am still not figuring this out

OpenStudy (jkristia):

what is 'r' the radius?. If you have a triangle tan is b / a, where a is the 'x axis' leg and b is the 'y axis' leg

OpenStudy (jkristia):

you can find sin and cos with this information. You know r = ?

OpenStudy (anonymous):

Yes, r=sqrt10, this is confirmed in math book answer

OpenStudy (jkristia):

yes, so you have all the info to find sin and cos (and the rest)

OpenStudy (jkristia):

or it could be 3 and -1, either would work.

OpenStudy (anonymous):

If in trig function, tan= b/a, then which is a and which is b?

OpenStudy (anonymous):

if it can be either, than how can i know for example, what to put in sin, if sin = b/r

OpenStudy (anonymous):

or in cotangent, when cotangent = a/b

OpenStudy (anonymous):

how would i know what to put where in cotangent

OpenStudy (anonymous):

BTW jkristia, I realize i suck at this and am asking a lot of questions, I appreciate your help so much.

OpenStudy (jkristia):



OpenStudy (jkristia):

hmm, I couldnt see the image once I drew it. in this case the angle is the left angle (where r and a meet). In this case your trig functions refer to that one angle. so tan = b/a sin = b / r cos = a / r All related to that one angle. If you chose the top angle (where b and r meet) then your trigs are swapped tan = a / b, sin = a / r cos = b / r

OpenStudy (jkristia):

so if you know tan is -3, then you know that one of the leg is 3 (or - 3) and the other left is -1 (or 1), or any other values which gives the ratio -3

OpenStudy (anonymous):

OK, so essentially you are proving here that if I am given tan= -3, a and b are interchangeable.

OpenStudy (jkristia):

if you watch Khan's videos on trig, then he mention SOHKAHTOA many times. It means Sin = Opposite over Hypoternuse Cos = Adjacent over Hypoternuse Tan = Opposite over Adjacent

OpenStudy (jkristia):

no, the sign is, but not the values. The ratio of a/b must be -3.

OpenStudy (anonymous):

ok so for example, if I tried to answer this question, I could say that sin= 1/sqrt10 or sin=3/sqrt10, and either would be right?

OpenStudy (jkristia):

In your case you are refering to two different angles. You should pick one angle only. If you have tan = b / a = -3 and you chose b = -3 and a = 1, then you have sin = b / r = -3 / sqrt(10) cos = a / r = 1 / sqrt(10)

OpenStudy (anonymous):

oh sorry yes, sorry i am dislexic ... i used b=-3, a=1

OpenStudy (jkristia):

tan = b/a, then cot is a/b

OpenStudy (jkristia):

I think I'm mising the a's and the b's too. But if tan = -3, then cotan = 1/tan = -1/3

OpenStudy (anonymous):

Well thank you for all of your 30 minutes and effort.

OpenStudy (jkristia):

I had to restart IE, it gost stuck while writing an equation. \[\tan = \frac{b}{a}, \cot = \frac{a}{b} = \frac{1}{\tan}\] so what you have is correct.

OpenStudy (jkristia):

Did it get any clearer, or did I just manage to make you even more confused?. Just rememebr that the trig functions are for a given angle. If you pick another angle, the values will change.

OpenStudy (anonymous):

Well... with this limited information, can you just arbitrarily pick which one will be a and b, for this one example problem?

OpenStudy (jkristia):

no, tan is the opposite side over the adjacent side. That is the side opposite to the angle is 'b' and the side adjacent to the angle is 'a'. So now you know which is a and which is b, then you can find r, and all the other trig functions for that same angle. Maybe it would be good to watch the Khan videos on trigonometry, I found them very helpful. http://www.khanacademy.org/math/trigonometry

OpenStudy (anonymous):

because... and i promise im not being dyslexic this time... \[\tan= \frac{ b }{ a }, \] i used \[\tan(x)=\frac{ -3 }{ 1 }\]

OpenStudy (anonymous):

how in this problem am i supposed to figure out which one is adjacent and which one is opposite?

OpenStudy (jkristia):

Ahh - TAN is defined as opposite divided by adjacent. Since you have the -3/1 you already know that the side opposite to the angle is -3, and the side adjacent to the angle is 1 SOHCAHTOA SOH - sin = opposite over hyp \[\sin = \frac{-3}{\sqrt{10}}\] CAH - cos = adjacent over hyp \[\cos = \frac{1}{\sqrt{10}}\] TOA - tan = opposite over adjacent \[\tan = \frac{-3}{1}\]

OpenStudy (anonymous):

but it never tells where the angle is...

OpenStudy (jkristia):

found this link which shows the trigs http://www.mathwords.com/s/sohcahtoa.htm

OpenStudy (anonymous):

i watched the kahn academy video SOHCAHTOA 20 times

OpenStudy (anonymous):

btw thank you

OpenStudy (jkristia):

the angle is tan(x) = -3, so if you need to know the angle in degrees you need to take the arctan(-3) = 71.6 degress (if I got that right), but you dont need that to find the other trig values

OpenStudy (jkristia):

oops -71.6 of course. try tan(-71.6) and you get -3 then arctan (-3) and you get (-71.6). Anyway - I think I'm just making you more confused. Take a look the Khans videos, they are really good

OpenStudy (jkristia):

oh - sorry, missed you post regarding Khan.

OpenStudy (anonymous):

I just dont see how i am supposed to figure out which one is opposite and which one is adjacent without knowing where the angle starts off at

OpenStudy (jkristia):

maybe this will help, it has the triangle http://www.mathwords.com/s/sohcahtoa.htm

OpenStudy (jkristia):

I have to run, sorry I couldnt help

OpenStudy (anonymous):

thanks for your time.

OpenStudy (anonymous):

If I can weigh in... It does tell you where the angle is: in the second quadrant. Simply find the x and y values that would correspond to an angle that would appear in the second quadrant and have a value of -3.

OpenStudy (anonymous):

Ohhh.. woops.

OpenStudy (jkristia):

The angle could be in second and fourth quadrant, both could have a tan of -3

OpenStudy (anonymous):

but in his problem it states that the angle is in the second quadrant. So he should only focus on the second quadrant.

OpenStudy (jkristia):

>>side of x is in quadrant 2 argh - I missed that too

OpenStudy (jkristia):

so - I only made him more confused. I should have read the question properly first. Thanks

OpenStudy (anonymous):

OK. So in summary, because tan=-3 is in quadrant 2, it must be 3/-1 because tan=y/x, and in quadrant 2, the first coordinate (x and 1) is negative, and the second coordinate (y and 3) is positive, leaving me with only one solution, that tan=-3 is = to tan=3/-1

OpenStudy (jkristia):

correct.

OpenStudy (anonymous):

Thank you i am so happy i figured this out you have no idea haha

OpenStudy (jkristia):

and now you should be able to find the other trig values.